• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Z. Chai

    Log entry time 08:02:29 on June19,2000

    Entry number 44279

    keyword=Owl shift summary

    00:00 Start owl shift. Continue taking data at kin(135,180)
    Total charge 17.2C, need 0.4C more .

    00:28 MCC called to inform us some of their BPM monitors are not working

    00:30 The Main control of magnetic screen became dead. Closed all other
    X displays and Tried to restart it. The magnet screen couldn't be brought up
    because of some communication problem. Called Javier Gomez. He is going to
    check it from his home.

    1:10 Javier called back. And the main control is restarted successfully. But still
    couldn't check the HV from it.
    Replayed the recent runs, everything looks fine.
    1:15 MCC told us they fixed the BPM problem. The beam is stable now.

    1:41 finish run 2247. Put 6msr coll. in both Arms. Start taking pointing data for
    2:35 Start kinematics change to (90,180)
    Harp scan 931, coda 2253 junk, beam tripped)
    Harp scan 932, coda 2254 junk
    Harp scan 933, coda 2255 two peaks in the beam x, y distribution.
    Called MCC. MCC said it is because there are two lasers.

    Change carbon door to 3+6.
    MCC called back the beam position problem is fixed
    Harp scan 934 coda 2256 beam position good.

    3:00 Start to cycle the HQ2 and HQ3.
    HQ2 quenched. Asked MCC for an access to reset the HQ2.
    3:40 Adam and Rikki went down to the Hall to reset the HQ2.
    3:57 Ready to take data at new kin (90,180) .

    4:46 MCC told us they have problem with the IOC which controls the raster.
    Still trying to fix it.

    4:58 Beam back to Hall A. 5uA to check raster

    6:04 Take out colimators from both arms

    time run events charge comment
    23:51 2244 1.3M 0.1C
    00:17 2245 1.3M 0.1C
    00:43 2246 1.24M 0.1C
    01:17 2247 1.3M 0.1C Kin(135,180) finished. 17.55C.
    2248 junk
    02:02 2249 0.42M Pointing data 10cm dummy
    02:11 2250 15m dummy
    02:19 2251 4cm dummy
    02:28 2252 12C
    2255 harp 933 two peaks.
    2256 harp 934 harp scan. one peak
    2257 harp 935 optimizing beam position, worse than harp 934, two peaks, OTR in.
    2259 harp 937 good beam position now.

    4:07 2260 0.1M 12C New kin(-90) pointing data
    4:16 2261 0.07M 4cm dummy
    4:21 2262 0.2M 15cm dummy
    4:27 2263 0.2M 10cm dummy
    2264->2269 junk runs for test

    4:58 2270 short run to check spot size 4x4
    5:02 2271 1.2M ps1=ps2=ps3=ps43=1,ps5=65535, efficiency check
    5:07 2272 1.3M 146mC 6msr col. 76uA
    5:43 2273 0.5M 56mC 6msr col. 76uA
    6:03 2274 1.3M 87mC open col.
    6:27 2275 1.3M 87mC
    6:49 2276 1.3M 87mC
    7:09 2277 1.3M 87mC
    7:29 2278 1.3M 87mC
    7:52 2279